OpenStudy (anonymous):

Use a compound angle formula for cosine to show that cos2x=cos^2x-sin^2x

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

\[\cos(2x)=\cos(x+x)\]

OpenStudy (kc_kennylau):

Then apply the compound angle formula :)

OpenStudy (anonymous):

can you help me get it with the compound angle formula??

OpenStudy (anonymous):

It confuses me sooo much :(

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

What is confusing you? :)

OpenStudy (anonymous):

compound angle formula

OpenStudy (kc_kennylau):

But you just have to memorize the formulae by heart, I can't help you on this...

OpenStudy (anonymous):

but how do I use it??

OpenStudy (anonymous):

@kc_kennylau please heeeeeelp

OpenStudy (anonymous):

I suck at math :(

OpenStudy (kc_kennylau):

well, you use it when the thing inside the trig funcs are two things added together...

OpenStudy (kc_kennylau):

or one thing subtracting one thing

OpenStudy (anonymous):

can you do it for me.. so I can see hows it done??

OpenStudy (anonymous):

pleaseee

OpenStudy (kc_kennylau):

I'll do \(\sin(2x)\) for you so that I don't break the rules and you can learn :) \[\begin{array}{ll} &\sin(2x)\\ =&\sin(x+x)\\ =&\sin x\cos x+\cos x\sin x\\ =&\sin x\cos x+\sin x\cos x\\ =&2\sin x\cos x \end{array}\]

OpenStudy (anonymous):

Basically, this isn't technically cheating but you gotta see that when they have A and B, when you do cos(2x) its really just (x+x) and look at the addition formula, just replace all the As and Bs with xs

OpenStudy (anonymous):

so now I have to do it for sin?

OpenStudy (anonymous):

I mean cos! @kc_kennylau

OpenStudy (kc_kennylau):

yes :)

OpenStudy (anonymous):

okay check this @kc_kennylau

OpenStudy (anonymous):

cos(2x) cos(x+x) cosx sinx + cosx cosx cosx sin x + cosx sinx = 2cosx sinx I KNOW ITS WRONG

OpenStudy (anonymous):

thats the same link

OpenStudy (kc_kennylau):

The identity is \(\cos(A+B)=\cos A\cos B-\sin A\sin B\)

OpenStudy (kc_kennylau):

yep, it's always the same link

OpenStudy (kc_kennylau):

The formula for sine is different from the formula of cosine :)

OpenStudy (anonymous):

its wrong right?

OpenStudy (anonymous):

can you start it off for me then? :)

OpenStudy (kc_kennylau):

\[\cos(2x)\]\[=\cos(x+x)\]\[=\fbox{Now apply the formula :D}\]

OpenStudy (anonymous):

okay let me try it

OpenStudy (anonymous):

will it be 2cosx2sinx ?

OpenStudy (kc_kennylau):

nope... what do you get after applying the formula without simplifying? :)

OpenStudy (kc_kennylau):

Just plug in A=x and B=x into the formula \(\cos(A+B)=\cos A\cos B-\sin A\sin B\)...

OpenStudy (anonymous):

cosx cosx-sinx sinx

OpenStudy (kc_kennylau):

exactly :) now simplify it :)

OpenStudy (anonymous):

2sinxcosx

OpenStudy (anonymous):

??

OpenStudy (kc_kennylau):

how the hell did you get \(2\sin x\cos x\) from \(\cos x\cos x-\sin x\sin x\) /(-_-)\

OpenStudy (anonymous):

Idk

OpenStudy (kc_kennylau):

What do you get when you multiply a number by itself? :p

OpenStudy (anonymous):

square

OpenStudy (kc_kennylau):

exactly :D

OpenStudy (kc_kennylau):

so what do you get when you multiply \(\cos x\) by itself? :)

OpenStudy (anonymous):

cosx^2

OpenStudy (kc_kennylau):

and what do you get when you multiply \(\sin x\) by itself? :)

OpenStudy (anonymous):

sinx^2

OpenStudy (kc_kennylau):

and what about \(\cos x\cos x-\sin x\sin x\)?

OpenStudy (anonymous):

cosxsinx^2

OpenStudy (anonymous):

just tell me-_-

OpenStudy (kc_kennylau):

Side-note here, notation problem, we say \(\large\sin^2x\) instead of \(\large\sin x^2\) to avoid confusion, but they're just the same thing :) \[\cos(2x)\]\[=\cos(x+x)\]\[=\cos x\cos x-\sin x\sin x\]\[=(\cos x)^2-(\sin x)^2\]\[=\cos^x-\sin^2x\]

OpenStudy (anonymous):

thank you :)

OpenStudy (kc_kennylau):

wait, typo, should be \(=\cos^2x-\sin^2x\)